# Normalized mutual information is a biased measure for classification and community detection

### v1

@@ II.1 Variants of the mutual information
This answer is as wrong as it could be: we expect the mutual information to take the minimum value of zero, but instead it is equal to the entropy $H_{0}(g)$, which is the maximum possible value since $H_{0}(g)$ is an upper bound as we have said. In other words, $I_{0}(c;g)$ would have us believe that this algorithm which puts every object in its own group tells us everything there is to know about the true labeling $g$, when in fact it tells us nothing at all.
The reason for this result is that in this case the contingency table itself uniquely defines $g$, so neglecting it puts the mutual information in error by an amount equal to the complete information cost of the ground truth. If we include the cost of transmitting the contingency table on the other hand, this erroneous behavior disappears. In this case, we can calculate the number of contingency tables exactly:
$\Omega(n^{(c)},n^{(g)})=\frac{n!}{\prod_{s}n_{s}^{(g)}!},$ (22)
so the reduced mutual information is
$I(c;g)=I_{0}(c;g)-\log\frac{n!}{\prod_{s}n_{s}^{(g)}!}=0,$ (23)
which is now the expected answer.
A related shortcoming of the standard mutual information is that for finite $n$ a random labeling $c$ will have positive mutual information with respect to any ground truth in expectation: because mutual information is non-negative, fluctuations due to randomness in finite systems will produce non-negative values only and hence their average will in general be positive VEB10 ; Zhang15 . This seems counterintuitive: we would expect the average value for a random labeling to be zero.
We can rectify this issue by using another variant of the mutual information, which subtracts off the expected value of the measure, thereby making the average zero by definition. To do this we must first decide how a random labeling is to be defined—over what distribution of candidate labelings are we averaging? A popular choice is to take the uniform distribution over labelings that share the same group sizes $n^{(c)}$ as the actual candidate $c$. This yields the adjusted mutual information of Vinh et al. VEB10 :
@@
where the expectation is over the ensemble described. This gives $\langle I_{A}(c;g)\rangle_{\{c|n^{(c)}\}}=0$, so a positive value of the adjusted mutual information indicates a candidate labeling that performs above average within this ensemble.
However, the adjusted mutual information defined in this way, and particularly the choice of ensemble, has some problems NCY20 . The set of labelings with given $n^{(c)}$ heavily favors contingency tables with relatively uniform entries, simply because there are many more labelings that correspond to uniform tables than to non-uniform ones. Real contingency tables, on the other hand, tend to be highly non-uniform, since applications of the mutual information tend to focus on labelings that are at least somewhat similar to the ground truth (which produces a non-uniform table). This means that the average used in the adjusted mutual information puts most of its weight on configurations that are very different from those that occur in reality, making it a poor representation of the actual average mutual information.
A solution to this problem is to perform the average not over all labelings consistent with $n^{(c)}$ but over all contingency tables, which makes non-uniform contingency tables just as likely as uniform ones. As shown in NCY20 , however, adjusting the mutual information in this way simply gives us the reduced mutual information of Eq. (17) again. For this reason we favor the reduced mutual information over the adjusted mutual information: in addition to allowing for the information cost of the contingency table, it corrects sensibly for the nonzero expected value. It is also derived naturally from a fully information-theoretic argument, by contrast with the more ad hoc derivation of the adjusted mutual information. Notwithstanding, in Section III we give results using both reduced and adjusted mutual information, and find fairly similar outcomes in the two cases.
@@ II.2 Normalization of the mutual information
A fundamental difficulty with mutual information as a measure of similarity is that its range of values depends on the particular application, which makes it difficult to say when a value is large or small. Is a mutual information of 10 a large value? Sometimes it is and sometimes it isn’t, depending on the context. To get around this problem one commonly normalizes the mutual information so that it takes a maximum value of 1 when the candidate labeling agrees exactly with the ground truth. There are a number of ways this can be achieved but, as we show here, they are not all equal. In particular, some, including the most popularly used normalization, can result in biased results and should, in our opinion, be avoided. In its place, we propose a novel, unbiased normalized measure.
@@
$I_{0}^{(S)}(c;g)$ $=\frac{I_{0}(c;g)}{\frac{1}{2}[H_{0}(c)+H_{0}(g)]}$
$=\frac{I_{0}(c;g)}{\frac{1}{2}[I_{0}(c;c)+I_{0}(g;g)]}.$ (25)
This measure has a number of desirable features. Because of the inequalities in (13) its value falls between zero and one. And since both the base measure and the normalization are symmetric under interchange of $c$ and $g$, the normalized measure also retains this symmetry (hence the superscript “$(S)$,” for symmetric).
@@
Equation (25) is not the only normalization that achieves these goals however. Equation (13) implies that
@@
which gives us three more options for the denominator of the normalized measure. The arithmetic mean in Eq. (25), however, sees the most use by far LF09 ; OLC11 ; YAT16 ; fortunato2016community .
@@
We contend that in principle the unnormalized measure is more correct on this question, having a direct justification in terms of information theory. The purpose of the normalization is purely to map the values of the measure onto a convenient numerical interval, and should not change outcomes as it does here. Moreover, different symmetric normalizations can produce different results. For instance, if one normalizes by $\max(I_{0}(c;c),I_{0}(g;g))$ in Fig. 2 then candidate $c_{A}$ is favored in all cases.
@@
$I_{X}^{(A)}(c;g)=\frac{I_{X}(c;g)}{I_{X}(g;g)}.$ (30)
This definition still gives $I_{X}^{(A)}(g;g)=1$, but now the normalization factor in the denominator has no effect on choices between candidate labelings, since it is independent of $c$. In fact, Eq. (30) is the only way to normalize such that $I_{X}^{(A)}(g;g)=1$ while simultaneously ensuring that the preferred candidate is always the same as for the base measure. Thus this measure also removes any ambiguity about how one should perform the normalization.
The downside of this measure is the loss of the symmetry, which is undesirable in some applications. For instance, if one is comparing two candidate labelings directly to one another, rather than to a distinct ground truth, then a symmetric measure may be preferable. Embedding and visualization methods that employ mutual information as a similarity measure also normally demand symmetry PLC17 . In the most common applications of the normalized mutual information, however, where we are evaluating against a ground truth, which is an inherently asymmetric situation, the asymmetric measure makes perfect sense.
@@
$I^{(A)}(c;g)=\frac{I(c;g)}{I(g;g)},$ (31)
where
$I(c;g)=I_{0}(c;g)-\log\Omega(n^{(c)},n^{(g)}),$ (32)
with $\Omega(n^{(c)},n^{(g)})$ approximated as in Eq. (18). This measure correctly accounts for the information contained in the contingency table, returns value 0 in expectation when $c$ is uncorrelated with the ground truth and value 1 when $c=g$, and always favors the same labeling as the unnormalized measure.
One potential issue is that at present we are unable to prove a bound $I(c;g)\leq I(g;g)$ analogous to (13), which would ensure that $I^{(A)}(c;g)\leq 1$ always. There is strong numerical and other evidence that such a bound does apply—see Appendix B—but proving it is an open problem.
@@ III Example application:
As an example of the performance of our favored measure of mutual information and competing measures, we describe in this section the results of an extensive series of tests in which these measures are used to score the output of various algorithms for network community detection. In these tests we use the popular Lancichinetti-Fortunato-Radicchi (LFR) graph model LFR08 to generate random test networks with known community structure and realistic distributions of node degrees and group sizes, and then attempt to recover that structure using a variety of standard community detection algorithms, quantifying the accuracy of the recovery with six different measures: symmetrically normalized versions of the traditional mutual information, the adjusted mutual information, and the reduced mutual information, and asymmetrically normalized versions of the same. A number of studies have been performed in the past to test the efficacy of community detection algorithms on LFR benchmark networks LF09 ; OLC11 ; YAT16 ; fortunato2016community , but using only the symmetric non-reduced mutual information as a similarity measure. Our results indicate that this measure can produce biased outcomes and we recommend the asymmetric reduced mutual information instead.
@@
We perform community detection using six well-known algorithms as follows. (We use the implementations found in the igraph library CN06 , except for the inference method, for which we use the graph-tool library P14 .)
@@
    InfoMap: InfoMap is an information theoretic approach that defines a compression algorithm for encoding a random walk on a network based on which communities the walk passes through RB08 . Different community labelings give rise to more or less efficient compression, as quantified by the so-called map equation, and the labeling with the highest efficiency is considered the best community division.
@@
    Modularity maximization: Modularity is a quality function for network labelings equal to the fraction of edges within communities minus the expected such fraction if edge positions are randomized while preserving the node degrees. Modularity maximization aims to optimize the modularity over possible labelings and return the labeling with the highest modularity found. Modularity can be maximized using a number of heuristic methods, of which the most popular are agglomerative methods such as the Louvain and Leiden algorithms BGLL08 ; TWV19 , spectral methods Newman06b , and simulated annealing GSA04 ; MAD05 ; RB06a . In our tests we use simulated annealing where computationally feasible and the Leiden algorithm otherwise, these approaches giving the most consistent maximization of the modularity.
@@
    Modularity with enhanced resolution: Standard modularity maximization is known to suffer from a “resolution limit”—it cannot detect communities smaller than a certain threshold size FB07 . This can be remedied using a variant of modularity that includes a resolution parameter $\gamma$ such that higher values of $\gamma$ cause the algorithm to prefer smaller communities RB06a . Standard modularity maximization corresponds to $\gamma=1$, but for comparison we also conduct tests with $\gamma=10$ using the Leiden algorithm.
@@
    Statistical inference: Another popular approach to community detection makes use of model fitting and statistical inference. In this context the most commonly fitted model is the degree-corrected stochastic block model KN11a , which can be fitted using Bayesian methods to find the best community division Peixoto17 .
@@
    Walktrap: Walktrap is an agglomerative algorithm in which initially separate nodes are iteratively combined into progressively larger communities in order from strongest to weakest connections, where strength is quantified in terms of the time for a random walk to reach one node from another PL05 .
@@
    Labelprop: The label propagation or “labelprop” algorithm initially places every node in its own community then iteratively updates the labels of randomly chosen nodes by majority vote among their network neighbors, breaking ties at random RRS07 .
@@ III.1 Comparison between variants of the
Figure 3 summarizes the relative performance of the various mutual information measures in our tests. In this set of tests we limited ourselves, for the sake of clarity, to the top three community detection algorithms—InfoMap and the two variants of modularity maximization—and measured which of the three returned the best results according to each of our six mutual information measures, as a function of network size $n$ and the mixing parameter $\mu$. Each point in each of the six panels in the figure is color-coded with some mix of red, green, and blue to indicate in what fraction of cases each of the algorithms performs best according to the relevant measure and, as we can see, the results vary significantly among the measures. An experimenter trying to choose the best algorithm would come to substantially different conclusions depending on which measure they use.
@@
An extreme example of the disagreement between measures can be seen in the lower right corner of each plot in Fig. 3, where most measures return a negative value, indicating that community detection has failed altogether. This is expected: for very weak community structure all detection algorithms are expected to show a “detectability threshold” beyond which they are unable to identify any communities DKMZ11a ; Massoulie14 ; MNS15 . The standard normalized mutual information, however, claims to find community structure in this regime using the $\gamma=10$ version of modularity maximization. This occurs because the $\gamma=10$ algorithm finds many small communities and, as discussed in Section II.1, a labeling with many communities, even completely random ones, is accorded a high score by a non-reduced mutual information. This offers a clear reason to avoid the standard measure.
@@
The bottom left panel in Fig. 3 shows results for the asymmetrically normalized version of the traditional mutual information, which, perhaps surprisingly, gives even worse results than the symmetric version, with hardly any region in which the $\gamma=1$ version of modularity maximization outperforms the $\gamma=10$ version. The reason for this is interesting: the symmetric normalization of the traditional measure actually acts as an ad hoc correction against the errors introduced by neglecting the information content of the contingency table and hence reduces the amount by which the symmetric measure incorrectly rewards labelings with too many communities. The asymmetric normalization eliminates this correction and hence performs more poorly. The correct solution to this problem, however, is not to use a symmetric normalization, which can bias outcomes in other ways, but rather to adopt a reduced mutual information measure.
Finally, comparing the middle and right-hand columns of Fig. 3, we see that the results for the adjusted and reduced mutual information measures are quite similar in these tests, although there are some differences. In particular, the adjusted measure appears to find more significant structure for higher mixing than the reduced measure. This occurs because, as discussed in Section II.1 and Ref. NCY20 , the adjusted measure encodes the contingency table in a way that is optimized for more uniform tables than the reduced measure, and thus penalizes uniform tables less severely, leading to overestimates of the mutual information in the regime where detection fails. This provides further evidence in favor of using a reduced mutual information measure.
III.2 Comparison between community detection algorithms

### v2

@@ II.1 Reduced mutual information
This answer is not merely wrong; it is maximally so. The mutual information should take its minimum value of zero, but instead it takes the value $H_{0}(g)$, which is the maximum possible since $H_{0}(g)$ is an upper bound as we have said. The reason for this result is that in this case the contingency table itself uniquely defines $g$, so neglecting it puts the mutual information in error by an amount equal to the complete information cost of the ground truth. If we include the subleading terms on the other hand, this erroneous behavior disappears. Assuming for simplicity that the ground-truth groups are of equal size, the optimal concentration parameters are $\alpha_{g|c}=0$ and $\alpha_{g}=\infty$, and the reduced mutual information of Eq. (12) becomes
$I(c;g)$ $=I_{0}(c;g)-\log\left[\frac{n!}{\prod_{r}n_{r}!}(1/q_{g})^{n}\right]-n\log q_{g}$
$=0.$ (22)
which is the correct answer.
A related shortcoming of the standard mutual information is that for finite $n$ a random labeling $c$ will have positive mutual information with respect to any ground truth in expectation: because mutual information is non-negative, fluctuations due to randomness in finite systems will produce non-negative values only and hence their average will in general be positive [4, 9]. This seems counterintuitive: we would expect the average value for a random labeling to be zero.
We can rectify this issue by using another variant of the mutual information, which subtracts off the expected value of the measure, thereby making the average zero by definition. To do this we must first decide how a random labeling is to be defined—over what distribution of candidate labelings are we averaging? The conventional choice is to take the uniform distribution over labelings that share the same group sizes $n^{(c)}$ as the actual candidate $c$. This yields the adjusted mutual information of Vinh et al. [4]:
@@
where the expectation is over the relevant ensemble.
This measure can also be treated within the information-cost framework of the present paper, as described in [6]. There it is shown that the subtracted term $\langle I_{0}(c;g)\rangle_{\{c|n^{(c)}\}}$ is precisely equal to the average cost of transmitting the contingency table when labelings are drawn from the uniform distribution. Thus, one can view the adjusted mutual information as a reduced measure, but differing from the one proposed in this paper in two ways. First, the correction term counts the cost of transmitting the contingency table but not the cost of transmitting the counts $n^{(g)}$, as described in Section II. This, however, is only a minor concern—the latter is usually much smaller than the former and its neglect introduces only a small error in most cases.
The second difference is more serious and revolves around the choice to draw labelings uniformly from those with given $n^{(c)}$. As discussed in [6], this distribution heavily favors contingency tables with relatively uniform entries, simply because there are many more labelings that correspond to uniform tables than to non-uniform ones. Real contingency tables, on the other hand, are often highly non-uniform, since many applications of the mutual information focus on labelings that are somewhat similar to the ground truth (producing a non-uniform table). In such cases, the average used in the adjusted mutual information puts most of its weight on configurations that are very different from those that occur in reality, making it a poor representation of true information costs. The measure proposed in this paper, by contrast, deliberately allows for non-uniform tables by drawing them from a Dirichlet-multinomial distribution and we argue that this is a strong reason to favor it over the adjusted mutual information. In addition it is also derived naturally from a fully information-theoretic argument, by contrast with the more ad hoc derivation of the adjusted mutual information. Notwithstanding, in Section III we give results using both reduced and adjusted mutual information, and find fairly similar outcomes in the two cases.
@@ II.2 Normalization of the mutual information
A fundamental difficulty with mutual information as a measure of similarity is that its range of values depends on the particular application, which makes it difficult to say when a value is large or small. Is a mutual information of 10 a large value? Sometimes it is and sometimes it isn’t, depending on the context. To get around this problem one commonly normalizes the mutual information so that it takes a maximum value of 1 when the candidate labeling agrees exactly with the ground truth. There are a number of ways this can be achieved but, as we show here, they are not all equal. In particular, some, including the most popularly used normalization, can result in biased results and should, in our opinion, be avoided. In its place, we propose an alternative, unbiased normalized measure.
@@
$\text{NMI}_{0}^{(S)}(c;g)$ $=\frac{I_{0}(c;g)}{\frac{1}{2}[H_{0}(c)+H_{0}(g)]}$
$=\frac{I_{0}(c;g)+I_{0}(g;c)}{I_{0}(c;c)+I_{0}(g;g)}.$ (24)
This measure has a number of desirable features. Because of the inequalities in (16), its value falls strictly between zero and one. And since both the base measure and the normalization are symmetric under interchange of $c$ and $g$, the normalized measure also retains this symmetry (hence the superscript “$(S)$,” for symmetric).
@@
Equation (24) is not the only normalization that achieves these goals. Equation (16) implies that
@@
which gives us three more options for a symmetric denominator in the normalized measure. The arithmetic mean in Eq. (24), however, sees the most use by far [10, 11, 12, 13].
@@
We argue that the unnormalized measure is more correct on this question, having a direct justification in terms of information theory. The purpose of the normalization is purely to map the values of the measure onto a convenient numerical interval, and should not change outcomes as it does here. Moreover, different symmetric normalizations can produce different results. For instance, if one normalizes by $\max(I_{0}(c;c),I_{0}(g;g))$ in Fig. 2 then candidate $c_{A}$ is favored in all cases.
@@
$\text{NMI}_{X}^{(A)}(c;g)=\frac{I_{X}(c;g)}{I_{X}(g;g)}.$ (29)
This definition still gives $I_{X}^{(A)}(g;g)=1$, but now the normalization factor in the denominator has no effect on choices between candidate labelings, since it is independent of $c$. In fact, Eq. (29) is the only way to normalize such that $I_{X}^{(A)}(g;g)=1$ while simultaneously ensuring that the preferred candidate is always the same as for the base measure. Thus this measure also removes any ambiguity about how one should perform the normalization. Loosely, this asymmetrically normalized mutual information measures “how much $c$ tells us about $g$ as a fraction of all there is to know about $g$.”
The amount of bias inherent in the symmetrically normalized measure when compared with asymmetric one can be quantified by the ratio between the two:
${\text{NMI}_{X}^{(S)}(c;g)\over\text{NMI}_{X}^{(A)}(c;g)}$ $={[I_{X}(c;g)+I_{X}(g;c)]/[I_{X}(c;c)+I_{X}(g;g)]\over I_{X}(c;g)/I_{X}(g;g)}$
$=\frac{1+I_{X}(g;c)/I_{X}(c;g)}{1+I_{X}(c;c)/I_{X}(g;g)}.$ (30)
If the base measure $I_{X}$ is itself symmetric (which it is, either exactly or approximately, for all the measures we consider), then this simplifies further to
${\text{NMI}_{X}^{(S)}(c;g)\over\text{NMI}_{X}^{(A)}(c;g)}={H(g)\over\frac{1}{2}[H(g)+H(c)]}.$ (31)
Note how this expression results in a particular bias against complex candidate labelings $c$ that have higher entropy $H(c)$ than the ground truth. Traditional mutual information measures are particularly problematic when the candidate is a refinement of the ground truth, meaning the candidate groups are subsets of the ground-truth groups. The example given in Section II.1, in which the candidate labeling puts every node in a group on its own, is an extreme instance of this. Equation (31) shows that, to some extent, the traditional symmetric normalization corrects this issue: a candidate $c$ that takes the form of a refinement of $g$ will have $H(c)>H(g)$ and hence the normalized mutual information will be penalized.
We argue, however, that this is not the best way to address the problem and that the correct approach is instead to use the reduced mutual information. The reduced mutual information also penalizes such cases, but it does so in a more principled manner that directly addresses the root cause of the problem, rather than merely penalizing complex candidate labelings in an ad hoc manner as a side-effect of normalization. We will see examples of this in Section III, where we apply both measures to community detection in networks and find that, although the symmetrically normalized traditional measure is biased, it is in some cases fortuitously biased in the right direction, although it is still problematic in some others.
An obvious downside of asymmetric normalization is the loss of the symmetry in the final measure. In the most common applications of normalized mutual information, where labelings are evaluated against a ground truth, an inherently asymmetric situation, the asymmetric measure makes perfect sense, but in other cases the lack of symmetry can be undesirable. Embedding and visualization methods that employ mutual information as a similarity measure, for example, normally demand symmetry [14]. And in cases where one is comparing two candidate labelings directly to one another, rather than to a separate ground truth, then a symmetric measure may be preferable.
Even in this case, however, the asymmetric measure may sometimes be the better choice, as discussed in the introduction. For instance, when one labeling $c_{1}$ is a refinement of the other $c_{2}$, the information content is inherently asymmetric: $c_{1}$ says more about $c_{2}$ than $c_{2}$ does about $c_{1}$. An explicit example of this type of asymmetry is shown in Figure 3, where we consider two labelings of 27 objects. The left labeling $c_{1}$ is a detailed partition of the objects into nine small groups while the right labeling $c_{2}$ is a coarser partition into only three groups, each of which is an amalgamation of three of the smaller groups in $c_{1}$. Because of this nested relationship, it is relatively easy to transmit $c_{2}$ given knowledge of $c_{1}$ but more difficult to do the reverse. This imbalance is reflected in the asymmetrically normalized mutual information values in each direction (top and bottom arrows in the figure), but absent from the symmetric version (middle row).
@@
$\text{NMI}^{(A)}(c;g)=\frac{I(c;g)}{I(g;g)},$ (32)
where the mutual information $I(c;g)$ is quantified as in Eq. (12). This measure correctly accounts for the information contained in the contingency table, returns 0 in expectation when $c$ is uncorrelated with the ground truth and 1 when $c=g$, and always favors the same labeling as the unnormalized measure.
Figure 3: Values of the various normalized mutual information measures we consider, between two example partitions of the same set of 27 objects. The left partition $c_{1}$ divides the objects into nine groups of three objects each and is a refinement of the right partition $c_{2}$, which divides them into only three groups, aggregating those on the left. The arrows indicate the direction of the comparison and the accompanying notations give the values of the mutual information measures. All three asymmetrically normalized measures capture the intuition that the partition $c_{1}$ tells us more about $c_{2}$ than $c_{2}$ tells us about $c_{1}$.
@@ III Example application:
As an example of the performance of the various measures discussed above, we describe in this section the results of an extensive series of tests in which the measures are used to score the output of algorithms for network community detection. In these tests we use the popular Lancichinetti-Fortunato-Radicchi (LFR) graph model [15] to generate random test networks with known community structure and realistic distributions of node degrees and group sizes, and then attempt to recover that structure using a variety of standard community detection algorithms, quantifying the accuracy of the recovery with six different measures: symmetrically normalized versions of the traditional mutual information, the adjusted mutual information, and the reduced mutual information, and asymmetrically normalized versions of the same. A number of studies have been performed in the past to test the efficacy of community detection algorithms on LFR benchmark networks [10, 11, 12, 13], but using only the symmetrically normalized, non-reduced mutual information as a similarity measure. Our results indicate that this measure can produce biased outcomes and we recommend the asymmetric reduced mutual information instead.
@@
We perform community detection using six well-known algorithms as follows. (We use the implementations found in the igraph library [16], except for the inference method, for which we use the graph-tool library [17].)
@@
    InfoMap: InfoMap is an information theoretic approach that defines a compression algorithm for encoding a random walk on a network, based on which communities the walk passes through [18]. Different community labelings give rise to more or less efficient compression, as quantified by the so-called map equation, and the labeling with the highest efficiency is considered the best community division.
@@
    Modularity maximization: Modularity is a quality function for network labelings equal to the fraction of edges within communities minus the expected such fraction if edge positions are randomized while preserving the node degrees. The labeling with the highest modularity is considered the best community division. Modularity can be maximized using a number of heuristic methods, of which the most popular are agglomerative methods such as the Louvain and Leiden algorithms [19, 20], spectral methods [21], and simulated annealing [22, 23, 24]. In our tests we use simulated annealing where computationally feasible and the Leiden algorithm otherwise, these approaches giving the most consistent maximization of the modularity.
@@
    Modularity with a resolution parameter: Standard modularity maximization is known to suffer from a “resolution limit”—it cannot detect communities smaller than a certain threshold size [25]. This can be remedied using a variant of modularity that includes a resolution parameter $\gamma$ such that higher values of $\gamma$ cause the algorithm to prefer smaller communities [24]. Standard modularity maximization corresponds to $\gamma=1$, but for comparison we also conduct tests with $\gamma=10$ using the Leiden algorithm.
@@
    Statistical inference: Another popular approach to community detection makes use of model fitting and statistical inference. In this context the most commonly fitted model is the degree-corrected stochastic block model [26], which can be fitted using Bayesian methods to find the best community division [27].
@@
    Walktrap: Walktrap is an agglomerative algorithm in which initially separate nodes are iteratively combined into progressively larger communities in order from strongest to weakest connections, where strength is quantified in terms of the time for a random walk to reach one node from another [28].
@@
    Labelprop: The label propagation or “labelprop” algorithm initially places every node in its own community then iteratively updates the labels of randomly chosen nodes by majority vote among their network neighbors, breaking ties at random [29].
@@ III.1 Comparison between variants of the
Figure 4 summarizes the relative performance of the various mutual information measures in our tests. In this set of tests we limited ourselves, for the sake of clarity, to the top three community detection algorithms—InfoMap and the two variants of modularity maximization—and measured which of the three returned the best results according to each of our six mutual information measures, as a function of network size $n$ and the mixing parameter $\mu$. Each point in each of the six panels in the figure is color-coded with some mix of red, green, and blue to indicate in what fraction of cases each of the algorithms performs best according to each of the six measures and, as we can see, the results vary significantly among measures. An experimenter trying to choose the best algorithm would come to substantially different conclusions depending on which measure they use.
@@
An extreme example of the disagreement between measures can be seen in the lower right corner of each plot in Fig. 4, where most measures return a negative value, indicating that community detection has failed altogether. This is expected: for very weak community structure all detection algorithms are expected to show a “detectability threshold” beyond which they are unable to identify any communities [30, 31, 32]. The standard normalized mutual information, however, claims to find community structure in this regime using the $\gamma=10$ version of modularity maximization. This occurs because the $\gamma=10$ algorithm finds many small communities and, as discussed in Section II.1, a labeling with many communities, even completely random ones, is accorded a high score by a non-reduced mutual information. This offers a clear reason to avoid the standard measure.
@@
The bottom left panel in Fig. 4 shows results for the asymmetrically normalized version of the traditional mutual information, which gives even worse results than the symmetric version, with hardly any region in which the $\gamma=1$ version of modularity maximization outperforms the $\gamma=10$ version. This behavior arises for the reasons discussed in Section II.2: the bias inherent in the symmetric normalization fortuitously acts to partially correct the errors introduced by neglecting the information content of the contingency table. The asymmetric normalization eliminates this correction and hence performs more poorly. The correct solution to this problem, however, is not to use a symmetric normalization, which can bias outcomes in other ways as we have seen, but rather to adopt a reduced mutual information measure.
Finally, comparing the middle and right-hand columns of Fig. 4, we see that the results for the adjusted and reduced mutual information measures are quite similar in these tests, although there are some differences. In particular, the adjusted measure appears to find more significant structure for higher mixing than the reduced measure. This occurs because, as discussed in Section II.1 and Ref. [6], the adjusted measure encodes the contingency table in a way that is optimized for more uniform tables than the reduced measure, and thus penalizes uniform tables less severely, leading to overestimates of the mutual information in the regime where detection fails—in this regime the candidate and ground-truth labelings are uncorrelated which results in a uniform contingency table. This provides further evidence in favor of using a reduced mutual information measure.
III.2 Comparison between
community detection algorithms